find the mean of the probability distribution of the random variable x, which can take only the values if P(x)= 1/10 , for x= 1,2,3...., 10

1
Expert's answer
2022-03-03T17:25:17-0500

"\mu=\frac{1}{10}(1+2+3+4+5+6+7+8+9+10)=\frac{11}{2}=5.5."